a train whose speed is 75 mph, takes three minutes to pass a car that is going 60 mph in the same direction. How long is the train?
Answer by mananth(16949)

a train whose speed is 75 mph, takes three minutes to pass a car that is going 60 mph in the same direction. How long is the train?
..
difference in speed = 75-60 mph=15mph
time = 3minutes = 3/60 hours = 1/20 hours
length of train = 1/20 * 15
=3/4 mile
3/4 *5280=3960 feet
...
m.ananth@hotmail.ca